# Principles for Optimizing Quantum Transduction in Piezo-Optomechanical Systems

### v2

@@ IV.1.1 Constitutive equations of piezoelectricity
IV.1.2 Quantum Elastodynamics, and quantizing the Piezoelectric interaction
Here, $u_{i}(\vec{r})$ represents the $i^{th}$ component of the displacement of the medium from its original location given by vector $\vec{r}$ and dots over vector components (e.g., $\dot{u}_{i}$) represent time derivatives. In the approximation that we neglect rigid body rotation, we can replace the elements of the strain tensor $x_{ij}$ with corresponding spatial derivatives of displacement $du_{i}/dr_{j}$ (see Appendix D for details) [^3] The assumption of no rigid body rotation at the micro-scale means we will not be considering torsion waves in this system, which couple poorly to both the microwave and optical single modes in this system, but does not preclude them from being useful in other geometries.. This Lagrangian can be transformed to the Hamiltonian:
Here, the displacement can be decomposed as a sum over acoustic modes:
$u_{i}(\vec{r},t)=\sum_{n}\left(u_{n}^{(0)}w_{(in)}(\vec{r})e^{-i\omega_{n}t}+u_{n}^{(0)*}w_{in}^{*}(\vec{r})e^{i\omega_{n}t}\right)$ (32)
where $u_{n}^{(0)}$ is an overall amplitude factor scaling between the normalized mode functions $w_{in}(\vec{r})$ and the total displacement $u_{i}(\vec{r},t)$.
The mode functions themselves $w_{in}(\vec{r})$ can be found as solutions to the elastic equations of motion:
$-\omega_{n}^{2}\rho(\vec{r})w_{in}(\vec{r})=\sum_{jk\ell}\frac{\partial}{\partial r_{j}}c_{ijk\ell}\frac{\partial w_{kn}(\vec{r})}{\partial r_{\ell}}$ (33)
The normalization of the acoustic mode functions $w_{in}(\vec{r})$ can be set by defining an effective mode volume for each mechanical mode:
$V_{eff}^{mech(n)}=\Bigg{(}\int d^{3}r\bigg{(}\frac{\sum_{i}|w_{in}(\vec{r})|^{2}}{\int d^{3}r\sum_{i}|w_{in}(\vec{r})|^{2}}\bigg{)}^{2}\Bigg{)}^{-1}.$ (34)
In short, we take the fraction of $\sum_{i}|w_{in}(\vec{r})|^{2}$ divided by its integral over all space as a normalized density (integrating to unity). The average value of this density is equal to the integral of its square over all space. The effective mode volume $V_{eff}^{mech(n)}$ is the volume, which when multiplied by this average density gives unity. With the mechanical mode volume defined, the normalization condition for the acoustic mode functions is given by the convention:
$\int d^{3}r\sum_{i}|w_{in}(\vec{r})|^{2}=V_{eff}^{mech(n)}$ (35)
Note that our definition of the mechanical mode volume (34) is independent of scale changes to $w_{in}(\vec{r})$.
$\int d^{3}r\rho(\vec{r})\sum_{i}w_{im}^{*}(\vec{r})w_{in}(\vec{r})=m_{eff}^{(n)}\delta_{mn}$ (36)
Here, $m_{eff}^{(n)}$ is the effective mass of mode $n$ of the transducer. We can also use this to define an effective mass density $\rho_{eff}$ generally equal to the material mass density, such that $\rho_{eff}V_{eff}^{(n)}=m_{eff}^{(n)}$.
$H_{elas}$ $=\sum_{m}m_{eff}^{(m)}\omega_{m}^{2}(u_{m}^{(0)}u_{m}^{(0)*}+u_{m}^{(0)*}u_{m}^{(0)})$ (37)
When applying the quantization procedure, we can use the relations:
$u_{n}^{(0)}e^{-i\omega_{n}t}$ $\rightarrow\sqrt{\frac{\hbar}{2m_{eff}^{(n)}\omega_{n}}}\hat{b}_{n}(t)$ (38a)
$u_{n}^{(0)*}e^{i\omega_{n}t}$ $\rightarrow\sqrt{\frac{\hbar}{2m_{eff}^{(n)}\omega_{n}}}\hat{b}_{n}^{\dagger}(t)$ (38b)
to obtain the quantum elastic Hamiltonian:
with displacement operator:
$\hat{u}_{i}(\vec{r},t)\!=\!\sum_{n}\!\!\sqrt{\frac{\hbar}{2m_{eff}^{(n)}\omega_{n}}}\left(w_{in}(\vec{r})\hat{b}_{n}(t)\!+\!w_{in}^{*}(\vec{r})\hat{b}_{n}^{\dagger}(t)\right)$ (40)
where:
$\hat{x}_{jk}=\frac{\partial\hat{u}_{j}}{\partial r_{k}}$ (42)
$\hat{D}_{i}(\!\vec{r},t\!)\!=\!i\!\!\sum_{n}\!\sqrt{\!\frac{\hbar\omega_{n}}{2\eta_{eff,(n)}V_{eff}^{EM(n)}}}\!\left(\!\mathcal{E}_{in}\!(\!\vec{r})\hat{c}_{in}\!(\!t)\!-\!\mathcal{E}_{in}^{*}\!(\!\vec{r})\hat{c}_{in}^{\dagger}\!(\!t)\!\!\right)$ (43)
with mode functions $\mathcal{E}_{in}\!(\!\vec{r})$ normalized by
$\int d^{3}r\sum_{ij}\eta_{ij}(\vec{r})\mathcal{E}_{jm}(\vec{r})\mathcal{E}_{in}^{*}(\vec{r})=\delta_{mn}\eta_{eff,(n)}V_{eff}^{EM(n)}$ (44)
$V_{eff}^{EM(n)}=\frac{\left(\int d^{3}r\sum_{ij}\eta_{ij}(\vec{r})\mathcal{E}_{jn}(\vec{r})\mathcal{E}_{in}^{*}(\vec{r})\right)^{2}}{\int d^{3}r\left(\sum_{ij}\eta_{ij}(\vec{r})\mathcal{E}_{jn}(\vec{r})\mathcal{E}_{in}^{*}(\vec{r})\right)^{2}}$ (45)
Note here, that we separate out the factor of $i$ in the definition of the field modes $\mathcal{E}_{in}(\vec{r})$. As an example of the properties of the spatial mode function, we may use the quantization procedure in [28] to examine $\mathcal{E}_{in}(\vec{r})$ in the case of Hermite-Gaussian beam propagation to be:
$\mathcal{E}_{in}(\vec{r})\propto\frac{g_{\mu}(x,y)}{\sqrt{\mathscr{L}_{z}}}e^{ik(n)z}(\vec{\epsilon}_{kn})_{i}$ (46)
where $g_{\mu}(x,y)$ is a normalized square integrable wavefunction defining the Hermite-Gaussian wavefunction of order $\mu$; $\mathscr{L}_{z}$ is the quantization length, which disappears after integration, and $(\vec{\epsilon}_{kn})_{i}$ is the $i^{th}$ component of a unit vector whose direction is defined along the polarization of light indexed by momentum $k$ and polarization index $s$. From this, we see that $\mathcal{E}_{in}(\vec{r})$ gives the spatial dependence of the $\omega=\omega_{n}$ frequency component of the electromagnetic field. In addition, we express the electromagnetic hamiltonian in terms of the electric displacement field $\vec{D}$ in order to make the evolution of the quantum fields fully consistent with Maxwell’s equations [29].
@@ IV.1.2 Quantum Elastodynamics, and quantizing the Piezoelectric interaction
$=\sum_{mn}\!\hbar g_{ijk(mn)}\!\!\left(\!\hat{c}_{im}(t)\!+\!\hat{c}_{im}^{\dagger}(t)\!\!\right)\!\!\left(\!\hat{b}_{jn}(t)\!+\!\hat{b}_{jn}^{\dagger}(t)\!\!\right)$ (48)
@@
$\boxed{g_{ijk(mn)}\!\equiv\!i\frac{\sqrt{\frac{\omega_{m}}{\omega_{n}}}}{4V_{eff}^{(mn)}}\!\!\sqrt{\!\frac{h_{ijk}^{2}}{\eta_{eff}^{(m)}\rho_{eff}}}\!\!\int\!\!\!d^{3}r\mathcal{E}_{im}\!(\vec{r})\frac{\partial w_{jn}\!(\vec{r})}{\partial r_{k}}}$ (49)
and we have consolidated electromagnetic and acoustic mode volumes into an electromechanical mode volume: $V_{eff}^{(mn)}=\sqrt{V_{eff}^{EM(m)}V_{eff}^{mech(n)}}$.
@@
IV.1.3 Analysis of Piezoelectric coupling
@@
The bulk properties affecting the piezoelectric coupling include the piezoelectric tensor $h_{ijk}$ (more commonly tabulated as $e_{ijk}=\epsilon_{i\ell}h_{\ell jk}$ so that $h_{ijk}=\eta_{im}e_{mjk}$ [26]) as well as mass density of the material $\rho_{eff}^{(n)}$ and the effective dielectric constant of the material $\epsilon_{eff}^{(m)}$.
@@
In general, selecting a bulk material for optimum piezoelectric coupling can be done by considering the ratio $(e_{ijk})^{2}/(\epsilon_{eff}\rho_{eff})$ as a whole, rather than attempting to control terms individually. See Table 4 in Appendix A for comparisons of different materials.
@@
$g_{ijk(mn)}\!\propto\!\frac{1}{\sqrt{V_{eff}^{EM(m)}V_{eff}^{mech(n)}}}\!\!\int\!\!\!d^{3}r\mathcal{E}_{im}\!(\vec{r})\frac{\partial w_{jn}\!(\vec{r})}{\partial r_{k}}$ (51)
Scaling up or down the overall volume of the piezoelectric medium should not substantially effect $g_{ijk(mn)}$ because of how the mode volumes are contained in the normalization of $\mathcal{E}_{im}\!(\vec{r})$ and $w_{jn}\!(\vec{r})$. If we assume that doubling the volume of the medium also doubles the effective mode volumes, then doing so does not change $g_{ijk(mn)}$; the values of $\mathcal{E}_{im}\!(\vec{r})$ and $w_{jn}\!(\vec{r})$ would each increase by a factor of $\sqrt{2}$, which would cancel out the decrease by the terms outside the integral. However, it is worth noting that one cannot change the dimensions of the medium in isolation. Other parameters, such as the resonant frequencies, would change as well.
@@
As this integral can be interpreted as a inner product between two functions in position space, we can also consider maximizing its overlap in momentum space. From this we get the intuition that we can maximize the coupling when the wavelengths of the microwave light and the acoustic vibrations are equal. However, this comes as the cost of bandwidth. When microwave light is propagating through the piezoelectric medium with a well-defined momentum (say a narrowband beam moving through a single-mode waveguide, or a collimated beam in a large bulk system), then only those acoustic modes with a correspondingly narrow range of momentum will be significantly excited. That said, the dimensions of the nonlinear medium impose lower limits to the momentum bandwidth of the acoustic waves (via the uncertainty principle), so that a smaller piezoelectric medium will have a larger range of wavelengths over which its acoustic oscillation will be resonant with the microwave input, and may therefore couple to a broader range of microwave frequencies.
If we instead consider a microwave mode with a broad momentum bandwidth, it may be coupled to many acoustic modes, but not strongly. Depending on the desired performance of the transducer (such as maximum fidelity at the expense of efficiency) this could be a fruitful design choice.
IV.2 The Optomechanical coupling constant
@@
The strength of the photoelastic effect is parameterized by the photoelastic tensor $p_{ijk\ell}$, which relates the strain $x_{k\ell}$ to the (relative) change in the inverse permittivity $\eta_{ij}$:
@@
See Appendix C for details. Since the displacement and electric fields are directly proportional to one another in this regime, the photoelastic tensor is directly related to the amount of applied stress generated as a quadratic function of the electric field (i.e., electrostriction).
At this point, we should note that electrostriction can be regarded as a contributing factor to the third-order ($\chi^{(3)}$) nonlinear optical susceptibility because the intensity-dependent stress induced in the material via electrostriction produces a strain relative to its elasticity $c_{ijk\ell}$, which in turn alters the refractive index through its photoelasticity (and therefore contributes to the intensity-dependent refractive index, which is a $\chi^{(3)}$-phenomenon) That said, multiple other processes also contribute to this effect such as localized heating of the material due to absorption, changing the polarizeability of the material independent of inducing strain, and saturated atomic absorption, among others [30].
It’s also worth mentioning that the photoelastic interaction is the basis for the scattering of light by sound waves (i.e., Brillouin scattering), but that light can be converted into phonons by other means. Where acoustic phonons are described by quantizing the vibrations of the material that come from fluctuations in the strain $x_{ij}$, there are also optical phonons describing vibrational modes in the medium that do not contribute significantly to an overall strain. These optical phonon modes occur between atoms within a crystal unit cell, vibrating against each other such that the center of mass of each unit cell is approximately constant (which implies the spacing between unit cells and therefore the overall strain are approximately unchanging as well)  [^4] See [46] for basic model describing optical phonons in a 1D atomic lattice with a two-atom unit cell, and see [47] for a thorough discussion of coupled oscillators necessary to extend optical phonons to multi-atom multi-dimensional unit cells. (See Appendix B for detailed discussion of this point). Scattering of light off of optical phonon modes is known as Raman scattering [32], and does not contribute toward optimizing the transduction efficiency other than as a source of noise.
IV.2.1 The Optomechanical Interaction Hamiltonian
@@
Although other contributions are present, we can envelop these effects into an effective photoelasticity tensor $p_{ijk\ell}^{eff}$, in a similar way as boundary effects may result in an effective refractive index in an optical waveguide.
@@
where the magnetic field modes $\hat{B}_{i}$ are defined relative to the electric displacement field modes $\hat{D}_{i}$ in a manner consistent with Maxwell’s equations.
@@
$\hat{H}$ $=\sum_{im}\hbar\omega_{m}\hat{a}_{m}^{\dagger}(t)\hat{a}_{m}(t)$
$\qquad-\sum_{imn}\hbar\bar{G}_{mn}\;\hat{a}_{im}^{\dagger}(t)\hat{a}_{im}(t)\left(\hat{b}_{n}(t)+\hat{b}_{n}^{\dagger}(t)\right)$ (58)
@@
$\bar{G}_{mn}\!\!$ $=\!\!\sqrt{\frac{\hbar}{32\rho_{eff}^{(n)}V_{eff}^{mech(n)}\!\epsilon_{0}^{2}(\eta_{eff}^{(m)})^{2}(\!V_{eff}^{(m)EM}\!)^{2}\omega_{n}}}p_{ijk\ell}^{(mn),eff}$
×$\times\int\!\!d^{3}r\,\mathcal{E}_{im}(\vec{r})\mathcal{E}^{*}_{jm}(\vec{r})\frac{\partial w_{kn}(\vec{r})}{\partial r_{\ell}}$ (59)
is the single-photon optomechanical coupling constant. This optomechanical interaction hamiltonian facilitates frequency shifts related to the value of $\bar{G}_{mn}(\hat{b}_{n}+\hat{b}_{n}^{\dagger})$ between the frequency modes in question. Note here, that because we are considering a single electromagnetic field mode being coupled to the acoustic mode, we do not include the summations over both multiple electromagnetic field modes.
In the transduction scheme considered here, there is a classically bright coherent state pump field in the telecom band driving conversion of GHz-scale phonons into telecom scale photons. Since the MRRs are being pumped in part by a classically bright driving laser, we may describe the optical field modes as a sum of a classical amplitude and a quantum fluctuation (e.g.,$\hat{a}\rightarrow\bar{a}+\delta\hat{a}$) and linearize the interaction hamiltonian, using the linearization approximation:
@@
discussed previously, which yields the linearized optomechanical interaction hamiltonian:
@@
IV.2.2 Analysis of optomechanical coupling

### v4

@@ III.1.1 Constitutive equations of piezoelectricity
Here we have neglected the effect of temperature changes to the electric field and to the stress, which would be modeled with pyroelectric and thermal expansion tensors (where strain is converted to stress via the elasticity tensor). The elasticity tensor here is defined at a constant electric displacement field. This is known as the piezoelectrically stiffened elasticity because the dipole moment built up in response to the applied stress works against it to bring the system back toward equilibrium.
With the constitutive equations of piezoelectricity, and a quantization of linear elasticity (also covered in [16] and in Appendix C), one can derive a quantum-mechanical Hamiltonian for the piezoelectric interaction in terms of the quantized electric displacement field $\hat{D}_{i}$ and quantized mechanical displacement operator $\hat{u}_{i}$, which are in terms of the microwave photon and GHz phonon creation/annihilation operators $\hat{c}_{(i,m)}$ and $\hat{b}_{(j,n)}$, respectively:
@@
$=\sum_{mn}\!\hbar g_{ijk(mn)}\!\!\left(\!\hat{c}_{(\!i,m\!)}(\!t\!)\!+\!\hat{c}_{(\!i,m\!)}^{\dagger}(\!t\!)\!\!\right)\!\!\left(\!\hat{b}_{(\!j,n\!)}(\!t\!)\!+\!\hat{b}_{(\!j,n\!)}^{\dagger}(\!t\!)\!\!\right)$ (9)
@@
$\boxed{g_{ijk(mn)}\!\equiv\!i\frac{\sqrt{\!\frac{\omega_{m}}{\omega_{n}}}}{4V_{\!eff}^{(mn)}}\!\!\sqrt{\!\frac{h_{ijk}^{2}}{\eta_{\text{eff}}^{(m)}\rho_{\text{eff}}}}\!\!\int\!\!\!d^{3}r\mathcal{E}_{(i,m)}\!(\vec{r})\frac{\partial w_{(j,n)}\!(\vec{r})}{\partial r_{k}}}$ (10)
Note that this expression subtly differs from that in [16] due to our quantizing the electromagnetic field in terms of electric displacement field operators, which is needed to make the dynamics consistent with Maxwell’s equations [25]. In this expression, we have that: $\omega_{m}$ and $\omega_{n}$ are the frequencies of the microwave and mechanical modes of order $m$ and $n$ respectively; $V_{\text{eff}}^{(mn)}$ is the effective electromechanical mode volume for microwave mode $m$ and mechanical mode $n$ (taken to be approximately the volume of the oscillating system); $\eta_{\text{eff}}^{(m)}$ is the effective inverse dielectric permittivity for microwave mode $m$ (approximated as the inverse dielectric permittivity of the material); $h_{ijk}$ is the piezoelectric tensor (in stress-voltage form); $\rho_{\text{eff}}$ is the effective density of the mechanical mode, which we approximate as the density of the mechanical oscillator; $\mathcal{E}_{(i,m)}(\vec{r})$ is the $i^{th}$ component of the microwave spatial mode function of order $m$; and $w_{(j,n)}(\vec{r})$ is the $j^{th}$ component of the mechanical displacement spatial mode function of order $n$, so that its derivative with respect to $r_{k}$ is an expression of the strain in this mechanical mode. That the strain tensor can be described as a single derivative of mechanical displacement works on the assumption that there is no rigid body rotation or torsion within the material (as shown in Appendix E).
Comparing the piezoelectric coupling tensor $g_{ijk(mn)}$ to our original hamiltonian for the total system and its simplified equations of motion, we have:
@@
where $\Gamma$ is the total microwave linewidth of the MRR, equal to the sum of the intrinsic microwave linewidth $\Gamma_{0}$, and the external microwave coupling rate $\Gamma_{ex}$. Where $g_{EM}=ig_{ijk(mn)}$, we obtain a real value for $\sqrt{\gamma_{ex}}$. In general, $\gamma_{ex}$ is a measured quantity rather than derived, but we now see how changing the piezoelectric coupling $g_{EM}$ affects it.
III.1.2 Analysis of Piezoelectric coupling
@@
The bulk properties affecting the piezoelectric coupling include the piezoelectric tensor $h_{ijk}$ (more commonly tabulated as $e_{ijk}=\epsilon_{i\ell}h_{\ell jk}$ so that $h_{ijk}=\eta_{im}e_{mjk}$ [24]) as well as mass density of the material $\rho_{\text{eff}}^{(n)}$ and the effective dielectric constant of the material $\epsilon_{\text{eff}}^{(m)}$ at microwave frequencies.
@@
In general, selecting a bulk material for optimum piezoelectric coupling can be done by considering the ratio $(e_{ijk})^{2}/(\epsilon_{\text{eff}}\rho_{\text{eff}})$ (alternatively, $h_{ijk}^{2}\epsilon_{\text{eff}}/\rho_{\text{eff}}$) as a whole, rather than attempting to control terms individually. See Table 4 in Appendix A for comparisons of different materials. Note here, that the dielectric constant is taken at the GHz-scale frequencies, which will generally be substantially larger than what is seen at optical frequencies. For example, Barium Titanate has a tabulated dielectric constant of approximately $2000$, but this does not correspond to a huge refractive index in the telecom band (of order $n=45$). Instead, it is understood that the dielectric constant generally decreases with frequency until leveling off at unity at frequencies so high that the material cannot respond fast enough to the incoming light (i.e., doesn’t have enough time to be fully polarized before the direction of the electric field changes).
@@
$g_{ijk(mn)}\!\propto\!\frac{h_{ijk}}{\sqrt{V_{\text{eff}}^{EM(m)}V_{\text{eff}}^{mech(n)}}}\!\!\int\!\!\!d^{3}r\mathcal{E}_{(i,m)}\!(\vec{r})\frac{\partial w_{(j,n)}\!(\vec{r})}{\partial r_{k}}$ (12)
Scaling up or down the overall volume of the piezoelectric medium should not substantially effect $g_{ijk(mn)}$ because of how the mode volumes are contained in the normalization of $\mathcal{E}_{(i,m)}\!(\vec{r})$ and $w_{(j,n)}\!(\vec{r})$. If we assume that doubling the volume of the medium also doubles the effective mode volumes, then doing so does not change $g_{ijk(mn)}$; the values of $\mathcal{E}_{(i,m)}\!(\vec{r})$ and $w_{(j,n)}\!(\vec{r})$ would each increase by a factor of $\sqrt{2}$, which would cancel out the decrease by the terms outside the integral. However, it is worth noting that one cannot change the dimensions of the medium in isolation. Other parameters, such as the resonant frequencies, would change as well.
@@
One way this zero overlap can occur is when the polarizations of the electric field $\hat{E}_{i}(\vec{r})$ and the strain-induced electric field $h_{ijk}x_{jk}$ are orthogonal. Whether or not this can be achieved will depend on how the crystal class determines the piezoelectric tensor. Among the tabulated materials in the appendix with a particularly strong piezoelectric effect (i.e.,AlN, $\text{BaTiO}_{\text{3}}$, KTP, LBO, and ZnO), these happen to fall into crystal classes where there are only five distinct nonzero elements to the piezoelectric tensor (though symmetries between elements vary between classes). For these particularly piezoelectric materials, the strain-induced (piezo)electric field would have components $E_{1}=h_{15}x_{13}$; $E_{2}=h_{15}x_{23}$ and $E_{3}=h_{31}x_{11}+h_{32}x_{22}+h_{33}x_{33}$. Although somewhat complicated, we can see that normal strains affect only the $z$ component of the generated electric field, and shear strains affect only the transverse components of the electric field. Conversely, only the transverse components of the electric field can generate shear strain, and only the $z$ component of the electric field can generate normal strain. When designing the piezoelectric component of the transducer, and how it will interface with the optomechanical elements, this geometry must be taken into account.
Independent of polarization, the overlap integral can be interpreted as an inner product between two functions in position space. For more intuition, we can also consider examining this overlap in momentum space. From this we get the intuition that we can maximize the coupling when the wavelengths of the microwave light and the acoustic vibrations are equal (relative to the momentum uncertainty imposed by the dimensions of the medium). That said, the dimensions of the nonlinear medium impose lower limits to the momentum bandwidth of the acoustic waves (via the uncertainty principle), so that a physically smaller medium will accommodate a larger range of wavelengths over which its acoustic oscillation will be resonant with the microwave input.
III.2 The Optomechanical interaction
@@
The strength of the photoelastic effect is parameterized by the photoelastic tensor $p_{ijk\ell}$, which relates the strain $x_{k\ell}$ to the change in the (relative) inverse permittivity $\epsilon_{0}\eta_{ij}$:
@@
See Appendix C.2 for details. Since the displacement and electric fields are directly proportional to one another in this regime, the photoelastic tensor is directly related to the amount of applied stress generated as a quadratic function of the electric field (i.e., electrostriction).
Although the photoelastic interaction is the basis for the scattering of light by sound waves (i.e., Brillouin scattering), it is worth mentioning that light can be converted into phonons by other means. Where acoustic phonons are described by quantizing the vibrations of the material that come from fluctuations in the strain $x_{ij}$, there are also optical phonons describing vibrational modes in the medium that do not contribute significantly to an overall strain. These optical phonon modes occur between atoms within a crystal unit cell, vibrating against each other such that the center of mass of each unit cell is approximately constant (which implies the spacing between unit cells and therefore the overall strain are approximately unchanging as well)  [^4] See [48] for basic model describing optical phonons in a 1D atomic lattice with a two-atom unit cell, and see [49] for a thorough discussion of coupled oscillators necessary to extend optical phonons to multi-atom multi-dimensional unit cells. (See Appendix D for a detailed discussion of this point). Scattering of light off of optical phonon modes is known as Raman scattering [27], and does not contribute toward optimizing the transduction efficiency other than as a source of noise.
III.2.1 The Optomechanical Interaction Hamiltonian
@@
Although other contributions are present, we can envelop these effects into an effective photoelasticity tensor $p_{ijk\ell}^{\text{eff}}$, in a similar way as boundary effects may result in an effective refractive index in an optical waveguide.
@@
where the magnetic field modes $\hat{B}_{i}$ are defined relative to the electric displacement field modes $\hat{D}_{i}$ in a manner consistent with Maxwell’s equations [25].
@@
$\hat{H}$ $=\sum_{im}\hbar\omega_{m}\hat{a}_{m}^{\dagger}(t)\hat{a}_{m}(t)$ (19)
$\qquad-\sum_{imn}\hbar\bar{G}_{mn}\;\hat{a}_{(i,m)}^{\dagger}(t)\hat{a}_{(i,m)}(t)\left(\hat{b}_{n}(t)+\hat{b}_{n}^{\dagger}(t)\right)$
@@
$\bar{G}_{mn}\!\!$ $=\!\!\sqrt{\frac{\hbar}{32\rho_{\text{eff}}^{(n)}V_{\text{eff}}^{mech(n)}\!\epsilon_{0}^{2}(\eta_{\text{eff}}^{(m)})^{2}(\!V_{\text{eff}}^{(m)EM}\!)^{2}\omega_{n}}}p_{ijk\ell}^{(mn),\text{eff}}$
×$\times\int\!\!d^{3}r\,\mathcal{E}_{(i,m)}(\vec{r})\mathcal{E}^{*}_{(j,m)}(\vec{r})\frac{\partial w_{(k,n)}(\vec{r})}{\partial r_{\ell}}$ (20)
is the single-photon optomechanical coupling constant. To our knowledge, this is the first expression for the optomechanical coupling in transduction fully worked out in the literature. Enfolded in the optomechanical coupling constant is the supplementary concern of acoustic impedance matching between the micro-ring resonator, and the bulk medium in which it is embedded. By quantizing the mechanical oscillations of the bulk and (first) MRR together as a single object, the corresponding normal modes automatically include any boundary-related effects between acoustic waves propagating into/out of the MRR.
This optomechanical interaction hamiltonian facilitates frequency shifts related to the value of $\bar{G}_{mn}(\hat{b}_{n}+\hat{b}_{n}^{\dagger})$ between the frequency modes in question. Note here, that because we are considering a single electromagnetic field mode being coupled to the acoustic mode, we do not include the summations over both multiple electromagnetic field modes.
In the transduction scheme considered here, there is a classically bright coherent state pump field in the telecom band driving conversion of GHz-scale phonons into telecom-scale photons. Since the MRRs are being pumped in part by a classically bright driving laser, we may describe the optical field modes as a sum of a classical amplitude and a quantum fluctuation (e.g.,$\hat{a}\rightarrow\bar{a}+\delta\hat{a}$) and linearize the interaction hamiltonian, using the linearization approximation:
@@
(discussed in Appendix B.3), which yields the linearized optomechanical interaction hamiltonian:
@@
III.2.2 Analysis of optomechanical coupling